# hard problem of consciousness

### v2

@@ 9 Conclusions: Why Nature Does Not Like Zombies
In this paper we formally unified lower order states and higher order meta-representations, dissolving the hard problem of consciousness. We described a multilayered formalism illustrating how biological self-organising systems become phenomenally conscious when they construct a 1ST order self. A human lacking a 1ST order self could not perform causal reasoning needed to adapt as humans evidently can [pearl2018].
As previous research pointed out [bennett2024a], the computer metaphor with the seminal distinction between hardware and software is a simplification. Rather, software is a state of hardware. Nature privileges efficiency over abstract simplification. Compare the vast quantities of both training data and energy required by a large language model, to the small quantities humans need to solve a problem. Biological systems are more efficient, because they are adaptive at a every level [levin2024, bennett2024c]. Hence, rather than trying to explain the mind in the abstract like software, we started at the level of the embodied organism.
One consequence of our approach is that it places the phenomenal quality of conscious experience before access consciousness. We have shown a consistent definition of access consciousness requires 1ST and 2ND order selves. Phenomenal consciousness arises first, access comes later. Unlike panpsychism, we don’t believe rocks are conscious. Only self-organising systems that need to adapt, motivated by valence, while keeping track of the self are conscious. Consciousness is a necessary adaptation, and a philosophical zombie is impossible.
@@
This work was supported by a Fundaçao para a Ciencia e a Tecnologia (FCT) grant PTDC/FER-FIL/4802/2020 ; and 2020-02773 CEECIND FCT to A.C.
Ethics Approval
This article does not contain any studies with human participants or animals performed by any of the authors.
Appendix
@@
(notation) $E$ with a subscript is the extension of the subscript212121e.g. $E_{l}$ is the extension of $l$..
@@
. A proxy $<$ is a binary relation on statements. In this paper we use only one proxy, called the weakness proxy, which compares the cardinality of a statement’s extension. For statements $l_{1},l_{2}$ we have $l_{1}<l_{2}$ iff $\lvert Z_{l_{1}}\rvert<\lvert Z_{l_{2}}\rvert$. Whenever we use $<$ to compare statements, we are referring to the aforementioned weakness proxy.
@@
Proposition

### v3

@@ 9 Conclusions: Why Nature Does Not Like Zombies
We have described a multilayered formalism illustrating how biological self-organising systems become phenomenally conscious when they construct a 1ST order self. A human lacking a 1ST order self could not perform causal reasoning needed to adapt as humans evidently can [pearl2018].
As previous research pointed out [bennett2024a], the computer metaphor with hardware and software is a simplification. Rather, software is a state of hardware. Nature privileges efficiency over abstract simplification. Compare the vast quantities of both training data and energy required by a large language model, to the small quantities humans need to solve a problem. Biological systems are more efficient, because they are adaptive at a every level [levin2024, bennett2024c]. Hence, rather than trying to explain the mind in the abstract like software, we started at the level of the embodied organism.
One consequence of our approach is that it places the phenomenal quality of conscious experience before access consciousness. We have shown a consistent definition of access consciousness requires 1ST and 2ND order selves. Phenomenal consciousness arises first, access comes later. Unlike panpsychism, we don’t believe rocks are conscious. Only self-organising systems that need to adapt, motivated by valence, while keeping track of the self are conscious. Consciousness is an adaptation, and a philosophical zombie is impossible.
@@
Supported by a Fundaçao para a Ciencia e a Tecnologia (FCT) grant PTDC/FER-FIL/4802/2020; and 2020-02773 CEECIND FCT to A.C.
Appendix A Appendix
@@
(notation) $E$ with a subscript is the extension of the subscript181818e.g. $E_{l}$ is the extension of $l$..
@@
A proxy $<$ is a binary relation on statements. In this paper we use only one proxy, called the weakness proxy, which compares the cardinality of a statement’s extension. For statements $l_{1},l_{2}$ we have $l_{1}<l_{2}$ iff $\lvert Z_{l_{1}}\rvert<\lvert Z_{l_{2}}\rvert$. Whenever we use $<$ to compare statements, we are referring to the aforementioned weakness proxy.
@@
A.1 Proofs
Versions of or alternatives to proofs 1-3 were published in [bennett2023b]. Further details on these updated versions are available in on GitHub, along with the supporting experiments [bennett2024appendices].
Theorem 1 (sufficiency).
Assume $\alpha\sqsubset\omega$. The weakness proxy sufficient to maximise the probability that a parent $\omega$ is learned from a child $\alpha$ [^20] Assume there exist correct policies for $\omega$, because otherwise there would be no point in trying to learn it..
Proof.
You’re given the definition of $\mathfrak{v}$-task $\alpha$ from which you infer a hypothesis $\mathbf{h}\in{\Pi}_{\alpha}$. To learn $\omega$, you need $\mathbf{h}\in\Pi_{\omega}$:
  1. 1.
    For every $\mathbf{h}\in\Pi_{\alpha}$ there exists a $\mathfrak{v}$-task $\gamma_{\mathbf{h}}\in\Gamma_{\mathfrak{v}}$ s.t. $O_{\gamma_{\mathbf{h}}}=E_{\mathbf{h}}$, meaning $\mathbf{h}$ permits only correct outputs for that task regardless of input. We’ll call the highest level task $\gamma_{\mathbf{h}}$ s.t. $O_{\gamma_{\mathbf{h}}}=E_{\mathbf{h}}$ the policy task of $\mathbf{h}$ [^21] I’d like to give credit here to Nora Belrose for pointing out an error. Nora pointed out I was miscounting the number of tasks. As a result I realised I was not counting tasks, I was in fact counting policy tasks and had entirely neglected to mention this fact. This was a significant error which has now been corrected, with several additional steps added to account for equivalence..
  2. 2.
    $\omega$ is either the policy task of a policy in $\Pi_{\alpha}$, or a child thereof.
  3. 3.
    If a policy $\mathbf{h}$ is correct for a parent of $\omega$, then it is also correct for $\omega$. Hence we should choose $\mathbf{h}$ that has a policy task with the largest number of children. As tasks are uniformly distributed, that will maximise the probability that $\omega$ is $\gamma_{\mathbf{h}}$ or a child thereof.
  4. 4.
    For the purpose of this proof, we say one task is equivalent [^22] This is because switching from $\beta$ to $\zeta$ s.t. $I_{\beta}\neq I_{\zeta}$ and $O_{\beta}=O_{\zeta}$ would be to pursue the same goal in different circumstances. This is because inputs are subsets of outputs, so both sets of inputs are implied by the outputs. $O_{\zeta}$ implies $I_{\beta}$ and $O_{\beta}$ implies $I_{\zeta}$ to another if it has the same correct outputs.
  5. 5.
    No two policies in $\Pi_{\alpha}$ have the same policy task [^23] Every policy task for policies of $\alpha$ is non-equivalent from the others.. This is because all the policies in $\Pi_{\alpha}$ are derived from the same set inputs, $I_{\alpha}$.
  6. 6.
    The set of statements which might be outputs addressing inputs in ${I}_{\omega}$ and not ${I}_{\alpha}$, is $\overline{E_{{I}_{\alpha}}}=\{l\in L_{\mathfrak{v}}:l\notin E_{{I}_{\alpha}}\}$ [^24] This is because $E_{{I}_{\alpha}}$ contains every statement which is a correct output or an incorrect output, and $\overline{E_{{I}_{\alpha}}}$ contains every statement which could possibly be in $I_{\omega}$, $E_{I_{\omega}}$ and thus $O_{\omega}$..
  7. 7.
    For any given $\mathbf{h}\in{\Pi}_{\alpha}$, the extension $E_{\mathbf{h}}$ of $\mathbf{h}$ is the set of outputs $\mathbf{h}$ implies. The subset of $E_{\mathbf{h}}$ which fall outside the scope of what is required for the known task $\alpha$ is $\overline{E_{{I}_{\alpha}}}\cap E_{\mathbf{h}}$ [^25] This is because $E_{{I}_{\alpha}}$ is the set of all conceivable outputs by which one might attempt to complete $\alpha$, and so the set of all outputs that can’t be made when undertaking $\alpha$ is $\overline{E_{{I}_{\alpha}}}$ because those outputs occur given inputs that aren’t part of ${I}_{\alpha}$..
  8. 8.
    $\left|\overline{E_{{I}_{\alpha}}}\cap E_{\mathbf{h}}\right|$ increases monotonically with $\left|E_{\mathbf{h}}\right|$, because for all $e\in E_{\mathbf{h}}$ is $e\notin\overline{E_{{I}_{\alpha}}}$ iff $e\in E_{{I}_{\alpha}}$.
  9. 9.
    $2^{\left|\overline{E_{{I}_{\alpha}}}\cap E_{\mathbf{h}}\right|}$ is the number of non-equivalent parents of $\alpha$ to which $\mathbf{h}$ generalises. It increases monotonically with the weakness of $\mathbf{h}$.
  10. 10.
    Given $\mathfrak{v}$-tasks are uniformly distributed and $\Pi_{\alpha}\cap\Pi_{\omega}\neq\emptyset$, the probability that $\mathbf{h}\in{\Pi}_{\alpha}$ generalises to $\omega$ is
    $p(\mathbf{h}\in{\Pi}_{\omega}\mid\mathbf{h}\in{\Pi}_{\alpha},\alpha\sqsubset\omega)=\frac{2^{\left|\overline{E_{{I}_{\alpha}}}\cap E_{\mathbf{h}}\right|}}{2^{\left|\overline{E_{{I}_{\alpha}}}\right|}}$
$p(\mathbf{h}\in{\Pi}_{\omega}\mid\mathbf{h}\in{\Pi}_{\alpha},\alpha\sqsubset\omega)$ is maximised when $\left|E_{\mathbf{h}}\right|$ is maximised. Recall from definition 4 that $<_{w}$ is the weakness proxy. For statements $l_{1},l_{2}$ we have $l_{1}<_{w}l_{2}$ iff $\lvert E_{l_{1}}\rvert<\lvert E_{l_{2}}\rvert$. $\mathbf{h}$ that maximises $<_{w}$ will also maximise $p(\mathbf{h}\in{\Pi}_{\omega}\mid\mathbf{h}\in{\Pi}_{\alpha},\alpha\sqsubset\omega)$. Hence the weakness proxy maximises the probability that [^26] Subsequently it also maximises the sample efficiency with which a parent $\omega$ is learned from a child $\alpha$. a parent $\omega$ is learned from a child $\alpha$. $\square$
Theorem 2 (necessity).
To maximise the probability of learning $\omega$ from $\alpha$, it is necessary to use weakness as a proxy.
Proof.
Let $\alpha$ and $\omega$ be defined exactly as they were in the proof of prop. 1.
  1. 1.
    If $\mathbf{h}\in{\Pi}_{\alpha}$ and $E_{{I}_{\omega}}\cap E_{\mathbf{h}}={O}_{\omega}$, then it must be he case that ${O}_{\omega}\subseteq E_{\mathbf{h}}$.
  2. 2.
    If $\left|E_{\mathbf{h}}\right|<\left|{O}_{\omega}\right|$ then generalisation cannot occur, because that would mean that ${O}_{\omega}\not\subseteq E_{\mathbf{h}}$.
  3. 3.
    Therefore generalisation is only possible if $\left|E_{\mathbf{h}}\right|\geq\left|{O}_{\omega}\right|$, meaning a sufficiently weak hypothesis is necessary to generalise from child to parent.
  4. 4.
    For any two hypotheses $\mathbf{h}_{1}$ and $\mathbf{h}_{2}$, if $\mid E_{\mathbf{h}_{1}}\mid<\mid E_{\mathbf{h}_{2}}\mid$ then the probability $p(\left|E_{\mathbf{h}_{1}}\right|\geq\left|{O}_{\omega}\right|)<p(\left|E_{\mathbf{h}_{2}}\right|\geq\left|{O}_{\omega}\right|)$ because tasks are uniformly distributed.
  5. 5.
    Hence the probability that $\left|E_{{m}}\right|\geq\left|{O}_{\omega}\right|$ is maximised when $\left|E_{{m}}\right|$ is maximised. To maximise the probability of learning $\omega$ from $\alpha$, it is necessary to select the weakest hypothesis.
To select the weakest hypothesis, it is necessary to use the weakness proxy. $\square$
Theorem 3 (simplicity sub-optimality).
Description length is neither a necessary nor sufficient proxy for the purposes of maximising the probability that induction generalises.
Proof.
In propositions 1 and 2 we proved that weakness is a necessary and sufficient choice of proxy to maximise the probability of generalisation. It follows that either maximising $\frac{1}{\left|m\right|}$ (minimising description length) maximises $\left|E_{m}\right|$ (weakness), or minimisation of description length is unnecessary to maximise the probability of generalisation. Assume the former, and we’ll construct a counterexample with $\mathfrak{v}=\{a,b,c,d,e,f,g,h,j,k,z\}$ s.t. $L_{\mathfrak{v}}=\{\{a,b,c,d,j,k,z\},\{e,b,c,d,k\},\{a,f,c,d,j\},\{e,b,g,d,j,k,z\},\newline \{a,f,c,h,j,k\},\{e,f,g,h,j,k\}\}$ and a task $\alpha$ where
  • ${I}_{\alpha}=\{\{a,b\},\{e,b\}\}$
  • ${O}_{\alpha}=\{\{a,b,c,d,j,k,z\},\{e,b,g,d,j,k,z\}\}$
  • ${\Pi}_{\alpha}=\{\{z\},\{j,k\}\}$
Weakness as a proxy selects $\{j,k\}$, while description length as a proxy selects $\{z\}$. This demonstrates the minimising description length does not necessarily maximise weakness, and maximising weakness does not minimise description length. As weakness is necessary and sufficient to maximise the probability of generalisation, it follows that minimising description length is neither. $\square$
The following sketch is more for intuition and explanation than anything formal.
Theorem 4.